# Probing Solar Neutrino Deficit via Torsion-Induced Flavor Change in $f(T)$ Gravity

### v1

@@ Abstract
We investigate whether the spacetime torsion can modify neutrino flavor oscillations in $f(T)$ gravity. This offers a probe of modified teleparallel gravity in astrophysical environments. By using the Dirac action in teleparallel geometry, we derive an effective coupling between the torsion vector and neutrino current. In the weak-field limit around a spherical mass, we obtain analytical expressions for torsion-induced phase shifts and effective mass-squared differences. Our results indicate that both vacuum oscillations and the Mikheyev-Smirnov-Wolfenstein (MSW) resonance in matter are affected by these torsion-based modifications. Using solar neutrino data from Super-Kamiokande, SNO, Borexino, and KamLAND, we constrain the teleparallel model parameters and also the neutrino-torsion coupling.
@@ 1 Introduction
The discovery of neutrino flavor oscillation marked a major breakthrough in particle physics, providing the first direct evidence that neutrinos have mass. Proposed to solve the solar neutrino deficit [1, 2, 3, 4, 5, 7, 6], this phenomenon was confirmed by landmark experiments like Super-Kamiokande (SK) [8] and the Sudbury Neutrino Observatory (SNO) [9]. High-precision measurements of neutrino mass and mixing parameters have been achieved using a diverse array of sources, including solar [10, 11, 12], atmospheric [13, 14, 15], reactor [16, 17], and accelerator [18] neutrinos. The accepted theory is that neutrinos oscillate between three flavors as they travel. When passing through matter (like the Sun), the significant Mikheyev-Smirnov-Wolfenstein (MSW) effect [2, 19, 20] occurs, enhancing this flavor conversion. Consequently, neutrinos are probes of new physics, with fundamental questions about their mass origin and their Dirac or Majorana nature remaining pressing open issues.
Beyond the established phenomenon of neutrino oscillations in flat spacetime, their behavior in curved spacetime has become a significant area of research within extended gravity theories [6, 21, 4, 5, 22, 23, 24, 25, 26, 27]. Neutrinos play a crucial role in the early universe, and detecting the cosmic neutrino background (a goal of projects like PTOLEMY [28]) would powerfully test the standard cosmological model. Furthermore, neutrinos with varying masses have been proposed as a candidate to explain dark energy. In these models [31, 29, 30, 32, 33, 34, 35, 36], neutrinos and a scalar field are linked by the neutrino mass, which has roots in the dark energy model. The dependence of the neutrino mass on the environment due to the interactions with dark sectors demonstrates that the non-standard interactions between the scalar field and matter components [37, 38] modifies the MSW effect, and subsequently, generates damping signatures which triggers the idea that a second-order effect might contribute to the experimentally confirmed deficit of neutrino flux [31, 32, 33, 34].
@@
As mentioned above, this work establishes a direct link between quantum particle phenomenology and the geometric formulation of gravity. It provides a framework for testing $f(T)$ gravity model with current and future neutrino observatories such as JUNO [59], DUNE [60], and Hyper-Kamiokande [61]. If torsion effects are detected, they would offer new insights into the nature of dark energy, neutrino mass generation, and the unification of gravity with particle physics.
@@ 2 $f(T)$ Gravity and Spacetime Torsion
[…]
${\Gamma^{\lambda}}_{\mu\nu}=e^{\lambda}_{\hat{a}}~\partial_{\mu}e^{\hat{a}}_{\nu}$. (2)
@@
Under the local Lorentz group, the torsion tensor admits an irreducible decomposition into vector, axial, and purely tensorial components [40]. From the torsion tensor, one defines the torsion vector
@@
$A^{\mu}=\frac{1}{6}\epsilon^{\mu\alpha\beta\gamma}T_{\alpha\beta\gamma}.$ (7)
This quantity represents the totally antisymmetric part of the torsion tensor and is particularly relevant in theories with fermionic matter, where it naturally couples to axial currents. Here, $\epsilon^{\alpha\beta\gamma\mu}$ denotes the contravariant Levi–Civita tensor, which enables the dualization of antisymmetric tensors in curved spacetime. It is related to the Levi–Civita symbol $\tilde{\epsilon}^{\alpha\beta\gamma\mu}$—a tensor density of weight +1+1 with $\tilde{\epsilon}^{0123}=+1$—via the determinant of the metric g, i.e., we have
$\epsilon^{\alpha\beta\gamma\mu}=\frac{\tilde{\epsilon}^{\alpha\beta\gamma\mu}}{\sqrt{-\text{g}}},$ (8)
@@
that does not yield a constant torsion scalar. This tetrad eliminates inertial contributions and yields a consistent spherically symmetric solution in $f(T)$ gravity [68]. In this geometry, the axial torsion vector vanishes identically, so the axial couplings do not contribute. According to this tetrad field, the torsion scalar is explicitly given by
@@ 3 Torsion Effects on Neutrino Oscillations
This study investigates how spacetime torsion modifies neutrino flavor oscillations within teleparallel gravity, an alternative geometric formulation in which gravitation is mediated by torsion rather than curvature. Within the standard prescription, the torsion tensor $T^{\rho}{}_{\mu\nu}$ can be decomposed into three irreducible components: a vector part $T_{\mu}=T^{\lambda}{}_{\lambda\mu}$ (see Eq.(6)), an axial-vector part $A^{\mu}=\frac{1}{6}\epsilon^{\mu\nu\rho\sigma}T_{\nu\rho\sigma}$ (see Eq.(7)), and a purely tensorial part $t_{\lambda\mu\nu}$ [40, 77, 78]. This decomposition leads to an effective torsion-induced vector and axial-vector coupling Lagrangian, i.e., we have $\mathcal{L}^{\text{VA}}\propto\kappa^{V}T_{\mu}\,\bar{\psi}\gamma^{\mu}\psi+\kappa^{A}A_{\mu}\,\bar{\psi}\gamma^{\mu}\gamma^{5}\psi$, where $\kappa$’s denote the neutrino-torsion part coupling coefficients.
@@
Consequently, for the spherically symmetric spacetime examined in this work, the axial torsion vector $A^{\mu}$ vanishes for the adopted tetrad choice, leaving only the vector coupling contribution. Since the vector current identically vanishes for Majorana neutrinos, the subsequent analysis is therefore restricted only to the Dirac neutrinos.
@@ 3.1 Vacuum Neutrino Oscillation
can be interpreted as an effective potential induced by spacetime torsion. In fact, the quantity $\hat{E}_{\nu}$ is an effective energy appearing in the neutrino propagation equation after separating the spinor components. It is not the conserved Killing energy with the timelike Killing vector of the static metric. Instead, its spatial variation encodes the work done by the torsion on the neutrino as it propagates, analogous to a potential energy term in non-relativistic quantum mechanics. This torsion-induced contribution plays a role analogous to the matter potential in the MSW effect, but with a fundamentally geometric origin rather than arising from weak interactions with a background medium. The dependence on the component $T_{1}(x)$ indicates that neutrino propagation becomes sensitive to the local spacetime structure, leading to the position-dependent modifications of the neutrino oscillation phase.
@@
Here, $\Delta\kappa^{(ij)}=\kappa_{i}-\kappa_{j}$ represents the difference in how each mass eigenstate couples to the spacetime torsion field. This formulation demonstrates that the oscillation phase acquires a novel contribution from the integral of the torsion field along the neutrino’s path, in addition to the conventional phase-term dependent on the mass-squared difference $\Delta m_{ij}^{2}$.
@@ 3.2 Modified Matter Effects
These key oscillation parameters are naturally influenced by the Sun’s electron number density. Moving forward, it is simpler to treat neutrinos as if they were in a vacuum, but with re-defined mass and mixing parameters $\Delta m^{2}_{m}$ and $\theta_{m}$. It is evident that, even in vacuum, where the electron density $N_{e}$ approaches zero, the presence of torsion modifies the neutrino mass-squared splitting, i.e., $\Delta m^{2}_{m}\to\Delta\hat{m}^{2}\neq\Delta m^{2}$. For the allowed ranges of the model parameter $|\alpha|\lesssim 5~m^{2}$ (see the next section), the torsion-induced correction to $\Delta m_{31}^{2}$ is negligible compared to its vacuum value ($\sim 2.5\times 10^{-3}\text{eV}^{2}$ [84]), so, as previously mentioned, the two-flavor approximation remains valid for solar neutrinos.
@@ 4 Numerical Results
Figure 2: Radial and energy dependence of the effective solar mass-squared splitting. The mass-squared splitting tends to a constant value at large distances, corresponding to the LMA solution (cf. Table 1). This behavior reflects torsion-induced modifications to neutrino oscillations in vacuum. We have set the model parameter $\alpha\sim\mathcal{O}(1~m^{2})$ and two possible values for the coupling difference $|\Delta\kappa^{(ij)}|\sim\mathcal{O}(10^{-1})$ (left panel) and $|\Delta\kappa^{(ij)}|\sim\mathcal{O}(10^{-2})$ (right panel). Solid curves represent positive $\Delta\kappa^{(ij)}$ values, and dashed curves represent negative ones.
@@
Figure 3: Transition between matter and vacuum-dominated neutrino oscillation regimes. This figure is plotted for various values of the model parameter $\alpha$ (in units $m^{2}$). We have assumed that $\Delta\kappa\simeq 10^{-3}$. The observational data are taken from Borexino [76] (gray points) and SNO+SK [87] (a single black point).
To quantify the goodness-of-fit in probability space, the chi-square analysis is calculated using the formula
@@
Table 2 presents the constraints on the $f(T)$ gravity parameter $\alpha$ derived from measurements of the solar 8B neutrino flux reported by successive generations of solar neutrino experiments. For each dataset, the measured electron-neutrino flux $\Phi_{i}$ is used to define the corresponding observed survival probability, $P_{i}^{\mathrm{obs}}=\Phi_{i}/\Phi_{\mathrm{SSM}}$, where $\Phi_{\mathrm{SSM}}=(5.94\pm 0.65)\times 10^{6}\mathrm{cm^{-2}s^{-1}}$ denotes the 8B neutrino flux predicted by the Standard Solar Model (SSM) [BPS08(GS)] [88, 89]. The uncertainty $\sigma_{i}$ is evaluated by propagating both the experimental measurement errors and the uncertainty associated with the SSM normalization. These survival probabilities are then analyzed within the framework of the present model to obtain the best-fit values of $\alpha$, together with their corresponding $1\sigma$ and $2\sigma$ confidence intervals, for each individual experiment as well as for combined datasets.
The Kamiokande [90] result yields a best-fit value $\alpha_{\text{bf}}=0.22$ with relatively broad $1\sigma$ and especially $2\sigma$ confidence intervals, implying its limited statistical precision compared to later experiments. Therefore, Kamiokande provides only weak constraints on $\alpha$ and is fully consistent with $\alpha=0$ limit, i.e., the standard teleparallel gravity limit $f(T)=T$.
A general observation is that experiments with higher statistical precision, notably the SK phases I-IV [91, 92, 12, 8], yield comparatively tight constraints on $\alpha$. Individually, the SK phases favor positive best-fit values of $\alpha$ of order $\sim\mathcal{O}(1~m^{2})$, with best fits ranging from $\alpha_{\text{bf}}\simeq 0.55$ to $0.89$. Importantly, all individual SK measurements are statistically consistent with $\alpha=0$ at the $1\sigma$ level, as their confidence intervals include zero. The combined SK analysis (global-SK) significantly sharpens these constraints, reducing the allowed parameter space to [0.03,1.19][-0.03,1.19] at $1\sigma$ and [0.62,1.95][-0.62,1.95] at $2\sigma$. This highlights the strong constraining power of homogeneous, high-statistics datasets and suggests only mild deviations from the standard teleparallel gravity limit.
@@
The Borexino [76, 94] measurement, provides weaker constraints: its best-fit value is negative $\alpha_{\text{bf}}\simeq-1.07$, but both its $1\sigma$ and $2\sigma$ ranges include zero, rendering it consistent with both standard teleparallel gravity (with $\alpha=0$) and modest deviations (with $\alpha\neq 0$).
@@
In summary, Table 2 demonstrates that higher-precision solar neutrino measurements, particularly from Super-Kamiokande, place stringent bounds on $\alpha$, while channel-dependent effects in SNO introduce some tension and broaden the global parameter space. The combined analyses mildly favor positive values of $\alpha$, but the standard teleparallel gravity limit (with $\alpha=0$) remains statistically viable at the $2\sigma$ level.
@@
These effects stem from $\Delta\kappa$’s role in scaling the matter potential, with positive/negative values enhancing or suppression conventional flavor conversions. Current solar neutrino data constrain $|\Delta\kappa|\lesssim 2\times 10^{-3}$, while future experiments like JUNO [59] may impose tighter constraints.
@@
The SK phases I-IV exhibit remarkable internal consistency. Individually, all SK phases favor positive best-fit values of $\Delta\kappa_{\text{bf}}$ around $(1.5-1.7)\times 10^{-3}$, with overlapping confidence intervals. While the $2\sigma$ ranges of the individual phases extend into negative values, their $1\sigma$ intervals remain strictly positive, suggesting a mild but systematic preference for nonzero $\Delta\kappa$. This preference is significantly reinforced in the combined SK (global-SK) analysis, which yields $\Delta\kappa=(1.69^{+0.58}_{-0.53})\times 10^{-3}$ at $1\sigma$. Notably, the global-SK $2\sigma$ interval remains entirely positive, effectively implying $\Delta\kappa\neq 0$ within this dataset.
@@
The Borexino measurement yields a small best-fit value and extremely wide confidence intervals that span both positive and negative values over a broad range. Consequently, Borexino places only weak constraints on $\Delta\kappa$ and remains fully consistent with both the null hypothesis and the values favored by SK and SK+SNO.
@@
The dependence of the electron-neutrino survival probability $P_{ee}$ on the parameter $\alpha$ for $10$ MeV 8B neutrinos is shown in Figure 5(a), comparing with experimental results from several detectors (SNO, SK, Borexino, Kamiokande) and various global-fits (shown by the asterisks in the plot). As $\alpha$ increases, matter-induced effects encoded in the $f(T)$ framework become progressively suppressed, leading to a smooth rise in $P_{ee}$. In the regime $\alpha\gg 1~m^{2}$, the survival probability saturates at $P_{ee}\simeq 0.55$, indicating the recovery of the vacuum oscillation limit. The agreement with global solar neutrino data further supports the physical viability of the chosen parameter range.
Also, Figure 5(b) presents the dependence of the electron-neutrino survival probability $P_{ee}$ on the coupling difference $\Delta\kappa[10^{-3}]$. For very small coupling differences $\Delta\kappa\ll 10^{-3}$, $P_{ee}$ attains its minimum, indicating a regime dominated by standard matter effects. As $\Delta\kappa$ increases, deviations induced by the modified coupling become significant, leading to a systematic enhancement of $P_{ee}$. This trend reflects the growing influence of profile-dependent corrections in the neutrino propagation dynamics.
@@
By comparing different values of the parameters $\alpha$ (left panel) and $\Delta\kappa$ (right panel), we plot the survival probability $P_{ee}$ for KamLAND in Figure 6. The model assumes that the torsion effects caused by an average constant Earth density and is computed using the KamLAND best-fit mass and mixing parameters [86] (see Table 1), resulting in deviations in the standard MSW effect pattern. Based on the experimental solar data from Borexino [76] (data points in the plots), the most plausible constraints on these parameters are estimated to be $|\alpha|\lesssim 5.5~m^{2}$ and $|\Delta\kappa|\lesssim 0.006$. These bounds are consistent with the findings in Tables 2 and 3; any deviation beyond them conflicts with experimental evidence.
Figure 6: Electron-neutrino survival probability as a function of energy for (left) varying $\alpha[m^{2}]$ and (right) varying $\Delta\kappa$, for the KamLAND’s best-fit oscillation parameters.

### v2

# Constraints on $f(T)$ Gravity from Solar Neutrino Conversion

@@ Abstract
We report new constraints on the parameters of $f(T)$ teleparallel gravity, derived from its imprints on neutrino flavor oscillations. Our analysis reveals that the presence of space-time torsion can alter both vacuum oscillations and the matter-enhanced Mikheyev–Smirnov–Wolfenstein (MSW) resonance. By using data from solar neutrino experiments: Super-Kamiokande, SNO, Borexino, and KamLAND, we perform a combined analysis to place the first direct observational bounds on the neutrino-torsion coupling and $f(T)$ parameters. These results establish neutrino phenomenology as a novel astrophysical probe for testing fundamental modifications of gravity.
@@ 1 Introduction
The discovery of neutrino flavor oscillation marked a major breakthrough in particle physics, providing the first direct evidence that neutrinos have mass. Proposed to solve the solar neutrino deficit [1, 2, 3], this phenomenon was confirmed by landmark experiments like Super-Kamiokande (SK) [4] and the Sudbury Neutrino Observatory (SNO) [5]. High-precision measurements of neutrino mass and mixing parameters have been achieved using a diverse array of sources, including solar [6, 7, 8, 9], atmospheric [10, 11, 12], reactor [13, 14], and accelerator [15] neutrinos. The accepted theory is that neutrinos oscillate between three flavors as they travel. When passing through matter (like the Sun), the significant Mikheyev-SmirnovWolfenstein (MSW) effect [16, 17] occurs, enhancing this flavor conversion. Consequently, neutrinos are probes of new physics, with fundamental questions about their mass origin and their Dirac or Majorana nature remaining pressing open issues.
Beyond the established phenomenon of neutrino oscillations in flat spacetime, their behavior in curved spacetime has become a significant area of research within extended gravity theories [18, 19, 20, 21, 22, 23, 24]. Neutrinos play a crucial role in the early universe, and detecting the cosmic neutrino background (a goal of projects like PTOLEMY [25]) would powerfully test the standard cosmological model. Furthermore, neutrinos with varying masses have been proposed as a candidate to explain dark energy. In these models [28, 26, 27, 29, 30, 31, 32], neutrinos and a scalar field are linked by the neutrino mass, which has roots in the dark energy model. The dependence of the neutrino mass on the environment due to the interactions with dark sectors demonstrates that the non-standard interactions between the scalar field and matter components modifies the MSW effect, and subsequently, generates damping signatures which triggers the idea that a second-order effect might contribute to the experimentally confirmed deficit of neutrino flux [28, 29, 30, 31].
@@
This work establishes a direct link between quantum particle phenomenology and the geometric formulation of gravity. It provides a framework for testing $f(T)$ gravity model with current and future neutrino observatories such as JUNO, DUNE, and Hyper-Kamiokande. If torsion effects are detected, they would offer new insights into the nature of dark energy, neutrino mass generation, and the unification of gravity with particle physics.
@@ 2 $f(T)$ Gravity and Spacetime Torsion
where Greek indices label spacetime coordinates and Latin indices with a hat correspond to locally inertial Lorentz frames, with $\eta_{\hat{a}\hat{b}}=\mathrm{diag}(+1,-1,-1,-1)$. Within the framework of the TEGR, the Weitzenböck connection is constructed entirely from the tetrad fields and is given by
${\Gamma^{\lambda}}_{\mu\nu}=e^{\lambda}_{\hat{a}}~\partial_{\nu}e^{\hat{a}}_{\mu}$. (2)
@@
Under the Lorentz group, the torsion tensor admits an irreducible decomposition into vector, axial, and purely tensorial components [34]. From the torsion tensor, one defines the torsion vector
@@
$A^{\mu}=\epsilon^{\alpha\beta\gamma\mu}T_{\alpha\beta\gamma}.$ (7)
This quantity represents the totally antisymmetric part of the torsion tensor and is particularly relevant in theories with fermionic matter, where it naturally couples to axial currents. Here, $\epsilon^{\alpha\beta\gamma\mu}$ denotes the contravariant Levi–Civita tensor, which enables the dualization of antisymmetric tensors in curved spacetime. It is related to the Levi–Civita symbol $\delta^{\alpha\beta\gamma\mu}$—a tensor density of weight +1+1 with $\delta^{0123}=+1$—via the determinant of the metric g, i.e., we have
$\epsilon^{\alpha\beta\gamma\mu}=\frac{\delta^{\alpha\beta\gamma\mu}}{\sqrt{-\text{g}}},$ (8)
@@
that does not yield a constant torsion scalar. This tetrad eliminates inertial contributions and yields a consistent spherically symmetric solution in $f(T)$ gravity [61]. In this geometry, the axial torsion vector vanishes identically, so axial couplings do not contribute. According to this tetrad field, the torsion scalar is explicitly given by
@@ 3 Torsion Effects on Neutrino Oscillations
This study investigates how spacetime torsion modifies neutrino flavor oscillations within teleparallel gravity, an alternative geometric formulation in which gravitation is mediated by torsion rather than curvature. Within the standard prescription, the torsion tensor $T^{\rho}{}_{\mu\nu}$ can be decomposed into three irreducible components: a vector part $T_{\mu}=T^{\nu}{}_{\nu\mu}$ (see Eq.(6)), an axial-vector part $A^{\mu}=\frac{1}{6}\epsilon^{\mu\nu\rho\sigma}T_{\nu\rho\sigma}$ (see Eq.(7)), and a purely tensorial part $t_{\lambda\mu\nu}$ [34, 70, 71]. This decomposition leads to an effective torsion-induced vector and axial-vector coupling Lagrangian, i.e., we have $\mathcal{L}^{\text{VA}}\propto\kappa^{V}T_{\mu}\,\bar{\psi}\gamma^{\mu}\psi+\kappa^{A}A_{\mu}\,\bar{\psi}\gamma^{\mu}\gamma^{5}\psi$, where $\kappa$’s denote the neutrino-torsion part coupling coefficients.
@@
In the spherically symmetric geometry considered here, the axial torsion $A^{\mu}$ vanishes for the chosen tetrad leaving only the vector coupling term. For Majorana neutrinos, the vector current vanishes identically, so we restrict our analysis to Dirac neutrinos in what follows.
@@ 3.1 Vacuum Neutrino Oscillation
can be interpreted as an effective potential induced by spacetime torsion. The quantity $\hat{E}_{\nu}$ is an effective energy appearing in the neutrino propagation equation after separating the spinor components. It is not the conserved Killing energy with the timelike Killing vector of the static metric. Instead, its spatial variation encodes the work done by the torsion field on the neutrino as it propagates, analogous to a potential energy term in non-relativistic quantum mechanics. This torsion-induced contribution plays a role analogous to the matter potential in the MSW effect, but with a fundamentally geometric origin rather than arising from weak interactions with a background medium. The dependence on the component $T_{1}(x)$ indicates that neutrino propagation becomes sensitive to the local spacetime structure, leading to the position-dependent modifications of the neutrino oscillation phase.
@@
Here, $\Delta\kappa^{(ij)}=\kappa_{i}-\kappa_{j}$ represents the difference in how each mass eigenstate couples to the spacetime torsion field. This formulation demonstrates that the oscillation phase acquires a novel contribution from the integral of the torsion field along the neutrino’s path, in addition to the conventional phase-term dependent on the squared mass difference $\Delta m_{ij}^{2}$.
@@ 3.2 Modified Matter Effects
These key oscillation parameters are naturally influenced by the Sun’s electron number density. Moving forward, it is simpler to treat neutrinos as if they were in a vacuum, but with re-defined mass and mixing parameters $\Delta m^{2}_{m}$ and $\theta_{m}$. It is evident that, even in vacuum, where the electron density $N_{e}$ approaches zero, the presence of torsion modifies the neutrino mass-squared splitting, i.e., $\Delta m^{2}_{m}\to\Delta\hat{m}^{2}\neq\Delta m^{2}$. For the allowed ranges $|\alpha|\lesssim 5m^{2}$ the torsion-induced correction to $\Delta_{31}^{2}$ is negligible compared to its vacuum value ($\sim 2.5\times 10^{-3}eV^{2})$, so the two-flavor approximation remains valid for solar neutrinos.
@@ 4 Numerical Results
Figure 2: Radial and energy dependence of the effective solar mass-squared splitting. The mass-squared splitting tends toward the solar surface for higher-energy neutrinos, converging to a constant value, corresponding to the LMA solution (cf. Table 1). This behavior reflects torsion-induced modifications to neutrino oscillations in vacuum. We have set the model parameter $\alpha\sim\mathcal{O}(1~m^{2})$ and two possible values for the coupling difference $|\Delta\kappa^{(ij)}|\sim\mathcal{O}(10^{-1})$ (left panel) and $|\Delta\kappa^{(ij)}|\sim\mathcal{O}(10^{-2})$ (right panel). Solid curves represent positive $\Delta\kappa^{(ij)}$ values, and dashed curves represent negative ones.
@@
Figure 3: Transition between matter and vacuum-dominated neutrino oscillation regimes. This figure is plotted for various values of the model parameter $\alpha$ (in units $m^{2}$). We have assumed that $\Delta\kappa\simeq 0.001$. The observational data are taken from Borexino [69] (gray points) and SNO+SK [80] (a single black point).
To quantify the goodness-of-fit in probability space, the chi-square ($\chi^{2}$) analysis is calculated using the formula
@@
Table 2 summarizes constraints on the $f(T)$ gravity parameter $\alpha$ inferred from measurements of the solar 8B neutrino flux by several generations of solar neutrino experiments. The table reports, for each experiment or combination of experiments, the experimentally measured electron-neutrino flux and the corresponding best-fit value of $\alpha$, together with its $1\sigma$ and $2\sigma$ confidence intervals within the framework of the present model.
The Kamiokande [81] result yields a best-fit value $\alpha_{\text{bf}}=0.22$ with relatively broad $1\sigma$ and especially $2\sigma$ confidence intervals, implying its limited statistical precision compared to later experiments. Therefore, Kamiokande provides only weak constraints on $\alpha$ and is fully consistent with $\alpha=0$ limit.
A general observation is that experiments with higher statistical precision, notably the SK phases I-IV [82, 83, 84, 4], yield comparatively tight constraints on $\alpha$. Individually, the SK phases favor positive best-fit values of $\alpha$ of order $\mathcal{O}(1~m^{2})$, with best fits ranging from $\alpha_{\text{bf}}\simeq 0.55$ to $0.89$. Importantly, all individual SK measurements are statistically consistent with $\alpha=0$ at the $1\sigma$ level, as their confidence intervals include zero. The combined SK analysis (global-SK) significantly sharpens these constraints, reducing the allowed parameter space to [0.03,1.19][-0.03,1.19] at $1\sigma$ and [0.62,1.95][-0.62,1.95] at $2\sigma$. This highlights the strong constraining power of homogeneous, high-statistics datasets and suggests only mild deviations from the standard teleparallel gravity limit.
@@
The Borexino [69, 88] measurement, characterized by intermediate precision, provides weaker constraints: its best-fit value is negative $\alpha_{\text{bf}}\simeq-1.07$, but both its $1\sigma$ and $2\sigma$ ranges include zero, rendering it consistent with both standard teleparallel gravity (with $\alpha=0$) and modest deviations (with $\alpha\neq 0$).
@@
In summary, Table 2 demonstrates that higher-precision solar neutrino measurements, particularly from Super-Kamiokande, place stringent bounds on $\alpha$, while channel-dependent effects in SNO introduce some tension and broaden the global parameter space. The combined analyses mildly favor positive values of $\alpha$, but the standard teleparallel gravity limit remains statistically viable at the $2\sigma$ level.
@@
These effects stem from $\Delta\kappa$’s role in scaling the matter potential, with positive/negative values enhancing or suppression conventional oscillations. Current solar neutrino data constrain $|\Delta\kappa|\lesssim 2\times 10^{-3}$, while future experiments like JUNO [89] may impose tighter constraints.
@@
The SK phases I-IV exhibit remarkable internal consistency. Individually, all SK phases favor positive best-fit values of $\Delta\kappa_{\text{bf}}$ around $(1.5-1.7)\times 10^{-3}$, with overlapping confidence intervals. While the $2\sigma$ ranges of the individual phases extend into negative values, their $1\sigma$ intervals remain strictly positive, suggesting a mild but systematic preference for nonzero $\Delta\kappa$. This preference is significantly reinforced in the combined SK analysis, which yields $\Delta\kappa=(1.69^{+0.58}_{-0.53})\times 10^{-3}$ at $1\sigma$. Notably, the global-SK $2\sigma$ interval remains entirely positive, effectively implying $\Delta\kappa\neq 0$ within this dataset.
@@
The Borexino measurement, characterized by larger experimental uncertainties, yields a small best-fit value and extremely wide confidence intervals that span both positive and negative values over a broad range. Consequently, Borexino places only weak constraints on $\Delta\kappa$ and remains fully consistent with both the null hypothesis and the values favored by SK and SK+SNO.
@@
Figure 5(a) illustrates the dependence of the electron-neutrino survival probability $P_{ee}$ on the parameter $\alpha$ for $10$ MeV 8B neutrinos, comparing with experimental results from several detectors (SNO, SK, Borexino, Kamiokande) and global-fits (shown by the asterisks in the plot). As $\alpha$ increases, matter-induced effects encoded in the $f(T)$ framework become progressively suppressed, leading to a smooth rise in $P_{ee}$. In the regime $\alpha\gg 1~m^{2}$, the survival probability saturates at $P_{ee}\simeq 0.55$, indicating the recovery of the vacuum oscillation limit. The agreement with global solar neutrino data further supports the physical viability of the chosen parameter range.
Figure 5(b) presents the dependence of the electron-neutrino survival probability $P_{ee}$ on the coupling difference $\Delta\kappa[10^{-3}]$. For very small coupling differences $\Delta\kappa\ll 1$, the survival probability attains its minimum, indicating a regime dominated by standard matter effects. As $\Delta\kappa$ increases, deviations induced by the modified coupling become significant, leading to a systematic enhancement of $P_{ee}$. This trend reflects the growing influence of profile-dependent corrections in the neutrino propagation dynamics.
@@
Figure 6 illustrates the survival probability $P_{ee}$, comparing different values of the parameters $\alpha$ (left panel) and $\Delta\kappa$ (right panel). The model assumes that the torsion effects caused by an average constant Earth density and is computed using the KamLAND best-fit mass and mixing parameters (see Table 1), resulting in deviations in the standard MSW effect pattern. Based on the experimental solar data from Borexino [69] (data points in the plots), the most plausible constraints on these parameters are estimated to be $|\alpha|\lesssim 5.5~m^{2}$ and $|\Delta\kappa|\lesssim 0.006$. These bounds are consistent with the findings in Tables 2 and 3; any deviation beyond them conflicts with experimental evidence.
Figure 6: Electron-neutrino survival probability as a function of energy for (left) varying $\alpha[m^{2}]$ and (right) varying $\Delta\kappa$, for the KamLAND’s best-fit oscillation parameters from Table 1.